# Cluster Catch Digraphs with the Nearest Neighbor Distance

### v1

@@ 3 $v_{temp}\leftarrow\arg\max_{v\in V(D)}\{sc(v)\}$;
      5 $\hat{S}\leftarrow\hat{S}\cup\{v_{temp}\}$;
      6 $D\leftarrow D^{sub}(V_{temp})$;
@@
Greedy Algorithm 2 A greedy algorithm finding an approximate MDS which is greedy in a score function $sc(v)$ at each iteration.
3.1.1 Cluster Catch Digraphs Using a KS-Type Statistic
@@
CCDs, introduced by Devinney [14], are an adaptation of CCCDs [43] for unsupervised clustering. Given an unlabeled data set $\mathcal{X}=\{x_{1},x_{2},...,x_{n}\}\subset\mathbb{R}^{d}$, where points are drawn from a mixture distribution, the objective is to determine the number of clusters and the optimal data partitioning. Unlike CCCDs, CCDs determine the radius of each covering ball using a Kolmogorov-Smirnov (KS)-type statistic, which measures the deviation of the local density around each point $x_{i}$ from a null model such as Complete Spatial Randomness (CSR) [38].
The KS-statistic for a point $x_{i}$ is defined as:
$T_{KS}(x_{i},r)=F_{rw}(x_{i},r)-F_{0}(x_{i},r),$
@@
where $F_{rw}(x_{i},r)$ represents the number of points within the covering ball $B(x_{i},r)$ centered at $x_{i}$ with radius $r$, and $F_{0}(x_{i},r)$ is the expected number of points under the null model, typically given by $F_{0}(x_{i},r)=\delta r^{d}$ for CSR, where $\delta$ is a density parameter and $d$ is the dimension of the space. The radius $r_{x_{i}}$ is chosen to maximize $T_{KS}(x_{i},r)$:
$r_{x_{i}}=\arg\max_{r\geq 0}\{T_{KS}(x_{i},r)\},$
@@
ensuring that the points covered by $B(x_{i},r_{x_{i}})$ form a significant local cluster.
@@
Once the optimal radii $r_{x_{i}}$ are determined for all points, a CCD for $\mathcal{X}$ is constructed as a digraph $D=(V(D),A(D))$, where each vertex corresponds to a point in $\mathcal{X}$, and there is an arc from $x_{i}$ to $x_{j}$ if $x_{j}\in B(x_{i},r_{x_{i}})$. The weakly connected components of $D$ are taken as clusters.
@@
To reduce the complexity of the covering, we apply the Greedy Algorithm 1 to find an approximate MDS, denoted $\hat{S}$, which selects a subset of covering balls to represent the clusters more efficiently [38]. The algorithm iteratively selects the vertex with the largest outdegree and removes its neighborhood until all points are covered.
@@
To further reduce the complexity of the cluster cover, an intersection graph $G_{MD}=(V_{MD},E_{MD})$ is constructed, where $V_{MD}=\hat{S}$, and an edge exists between $u,v\in\hat{S}$ if their covering balls cover common points, i.e., $N(u)\cap N(v)\neq\emptyset$. The Greedy Algorithm 2 is then applied to $G_{MD}$ to prune the set $\hat{S}$ further, yielding a reduced set $\hat{S}(G_{MD})$ that represents the core clusters.
Finally, to refine the clustering further and remove noise or outlier clusters, the silhouette index $sil(x_{i})$ is used [23]. Clusters are added incrementally to the partition $P=\{P_{1},P_{2},...,P_{k}\}$ until the average silhouette index $sil(P)$ is maximized, ensuring only significant clusters are retained. The covering balls of the significant clusters are called dominating covering balls. Points not covered by any cluster are assigned based on the relative similarity measure $\rho(x_{i},B(x_{j},r_{x_{j}}))$:
$\rho(x_{i},B(x_{j},r_{x_{j}}))=\frac{d(x_{i},x_{j})}{r_{x_{j}}},$
where $d(x_{i},x_{j})$ is the distance between $x_{i}$ and the center $x_{j}$ of the nearest cluster.
3.1.2 Cluster Catch Digraphs using Ripley’s K Function
@@
While the silhouette index improves the robustness of KS-CCDs, some limitations remain, such as the inability to capture spatial distributions and the need for an unknown input density parameter $\delta$ [37]. To address this, Manukyan and Ceyhan [37] replaced the KS-type statistic with the Ripley’s K function [48], denoted as $K(t)$, to determine whether the points within each covering ball follow a homogeneous Poisson process (HPP). This test is referred to as the Spatial Randomness Monte Carlo Test (SR-MCT) with Ripley’s K function.
@@
The RK-CCD method sets the radius of each covering ball to the largest value for which the points inside the ball pass the SR-MCT, ensuring that the points encapsulated conform to a CSR pattern. The primary difference between RK-CCDs and KS-CCDs is in the determination of the radii, which is now based on Ripley’s K function rather than the KS-type statistic. Subsequent steps, such as reducing cover complexity using an approximate MDS and refining clusters with the silhouette index, remain the same as in KS-CCDs.
@@
A variant of RK-CCDs introduced by Manukyan and Ceyhan [37] allows for the detection of clusters with arbitrary shapes by considering the connected components of the intersection graph $G_{MD}$ as the final clusters instead of using the MDS.
@@ 4.1 The Limitations of RK-CCDs for Clustering
The Monte Carlo analysis in [53] reveals that outlier detection methods using U-MCCDs and SU-MCCDs exhibit low True Negative Rates (TNRs) for regular points in high-dimensional data sets. This degradation is attributed to the limitations of RK-CCDs in high-dimensional spaces, where dominating covering balls are not sufficiently large for latent clusters.
@@
Specifically, RK-CCDs present two key limitations [53]: (1) The initial points in the SR-MCT are not random, as they are always the centers of the covering balls, leading to inaccurate testing results, especially in high-dimensional data sets; (2) In SR-MCT, the point-wise confidence band for $K(t)$ is constructed on fixed $t$ values [37], which is not appropriate for high-dimensional data sets, where the distances between any two points increase substantially.
4.2 The SR-MCT with the Nearest Neighbor Distance
@@
To avoid the limitations of RK-CCDs, we have introduced a new CCD approach in [53], which conduct SR-MCT with the Nearest Neighbor Distance (NND) instead of the Ripley’s K function.
@@
Given a data set $\mathcal{X}=\{x_{1},x_{2},...,x_{n}\}$, the mean NND of $\mathcal{X}$ is $\bar{\mathcal{d}}=\frac{\sum_{i=0}^{n}\mathcal{d}_{i}}{n}$, where $\mathcal{d}_{i}$ represents the distance of $x_{i}$ to its nearest neighbor. We aim to compares $\bar{\mathcal{d}}$ with the expected mean NND (denoted as $\mu_{\mathcal{d}}$) under the assumption of CSR. A normality test offers a convenient way to obtain P-values for $\bar{\mathcal{d}}$. Because both $\mu_{\mathcal{d}}$ and $\sigma_{\bar{\mathcal{d}}}$ (the standard derivation of $\bar{\mathcal{d}}$ under the assumption of CSR) can be derived [10, 11]. However, when the sample size $n$ is small, the distribution of $\bar{\mathcal{d}}$ is skewed, making the normality test unfeasible [5]. Fortunately, a Monte Carlo test proposed by Besag and Diggle [5] provides an alternative, where the P-values are determined by the empirical quantiles.
@@
Consequently, we have proposed another CCD approach based on the SR-MCT with NND. Additionally, we made several meaningful key enhancements to improve the robustness and reliability of the new CCDs [53]: (1) Multiple Comparison with Holm’s Correction: Both the median and mean NND are considered for multiple comparisons, with a Holm’s Step-Down Procedure for correction [61]; (2) Descending Radius Testing: Larger covering balls are preferred for high-dimensional data sets. Therefore, we offer the option to test the candidate values for the radius in descending order; (3) Lower-Tailed Test: The test is lower-tailed as the upper tail is less relevant when the point pattern is significantly “regular”. (4) Exclusion of Center Points: The center points are excluded from the test as they are not random.
@@
The new clustering approach based on CCDs and NND is called Uniformity- and Neighbor-based CCD (UN-CCD) clustering method, which is presented in Algorithm 3. It is worth noting that the only difference among KS-CCDs, RK-CCDs, and UN-CCDs is how to determine the radius of covering balls. The time and space complexities of UN-CCDs are $O((N+d+\log n)n^{2}+n^{3})$ and $O(n^{2})$ respectively [53].
Input: $\alpha$, data set $\mathcal{X}=\{x_{1},x_{2},...,x_{n}\}$;
Output: Cluster centers of $\mathcal{X}$;
Algorithm Steps:
1 foreach $x_{i}\in\mathcal{X}$ do
      2 Calculate distances $D(x_{i})=\{d(x_{i},x_{j})|x_{j}\in\mathcal{X},x_{i}\neq x_{j}\}$;
      3 foreach distance $r_{(j)}$ in $D(x_{i})$ sorted do
            4 Perform the SR-MCT with NND on $B(x_{i},r_{(j)})$;
            5 if test rejected at level $\alpha$ then
@@ 6 Set $r_{x_{i}}=r_{(j-1)}$; break;
7Construct a CCD $D=(V,A)$ using the pre-determined radii;
@@ 11 Output $\hat{S}(G_{MD})$ as cluster centers;
Algorithm 3 (UN-CCD Clustering Algorithm) Cluster Catch Digraphs based on the Nearest Neighbor Distance (NND). $\alpha$ is the level of the Monte Carlo test with NND.
@@
In Section 5, we compare the performance of KS-CCDs, RK-CCDs, and UN-CCDs with synthetic data sets consists of multiple clusters. We show the advantage of UN-CCDs over RK-CCDs on Gaussian clusters and high-dimensional space.
5 Monte Carlo Experiments for CCD-based Clustering
@@
Similar to the design of simulations in [53], the analysis begins with data sets consisting of uniform clusters.
@@
Additionally, the cluster centers are: (1) $1^{st}$ cluster: $\bm{\mu_{1}}=(\underbrace{3,...,3}_{d})$; (2) $2^{nd}$ cluster: $\bm{\mu_{2}}=(6,\underbrace{3,...,3}_{d-1})$; (3) $3^{rd}$ cluster: $\bm{\mu_{3}}=(6,6,\underbrace{3,...,3}_{d-2})$; (4) $4^{th}$ cluster: $\bm{\mu_{4}}=(3,5.5,\underbrace{3,...,3}_{d-2})$; (5) $5^{th}$ cluster: $\bm{\mu_{5}}=(8.5,\underbrace{3,...,3}_{d-1})$. The minimal distances between the first three cluster centers are 3, ensuring well separate clusters. On the other hand, the distances between the $1^{st}$ and the $4^{th}$, the $2^{nd}$ and the $5^{th}$ cluster centers are 2.5, leading in closer clusters, which may be confused as one cluster by many clustering methods. This study investigates the robustness of the three CCD-based clustering methods to clusters close in proximity. We present three realizations with different number of clusters under 2-dimensional space in Figure 2.
@@
Clustering quality is assessed using both internal and external validation measures, including the Adjusted Rand Index (ARI) [30], the average silhouette index (Sil) [49], and the frequency of successfully identifying the corrected number of clusters (denoted as success rate or % for simplicity). To mitigate the impact of randomness, we repeat each simulation 500 times, and the average measures across these repetitions are recorded.
RK-CCDs are “nearly” parameter-free, requiring only the significance level $\alpha$ of SR-MCT (with Ripley’s $K$ function). To accommodate for different dimensions, we set $\alpha$ to $1\%$ for $d<10$ and $0.1\%$ for $d\geq 10$, improving the performance of RK-CCDs. For UN-CCDs, the $\alpha$ of the SR-MCT (with NND) is set to $\{15\%,10\%,5\%,1\%,0.1\%\}$ as $d$ increases from 2 to 20. Unlike RK-CCDs and UN-CCDs, KS-CCDs require a density parameter $delta$ for the KS-based statistic. To manage the scale of $delta$ in high-dimensional settings, We work on the value of $d]{\delta}$ instead, and select the $d]{\delta}$ maximizing the average silhouette index in each simulation, presented in Table 2 and 6. The simulation results are summarized in Table 1, 3, and 4.
@@
Table 1: The simulation results of KS-CCDs on a set of synthetic data sets consisting of uniform clusters.
Density Parameter: $d]{\delta}$ $d=2$ $d=3$ $d=5$ $d=10$ $d=20$
$n=50$ 2 clusters 3.70 2.00 1.30 0.75 0.49
3 clusters 3.10 1.80 1.20 0.700 0.55
5 clusters 2.60 1.60 1.10 0.800 0.69
$n=100$ 2 clusters 5.20 2.50 1.50 0.80 0.53
3 clusters 4.40 2.40 1.30 0.78 0.56
5 clusters 3.80 2.10 1.30 0.88 0.70
$n=200$ 2 clusters 7.00 3.00 1.50 0.85 0.60
3 clusters 5.90 3.10 1.50 0.82 0.57
5 clusters 5.40 2.70 1.50 0.98 0.74
$n=500$ 2 clusters 11.00 3.90 1.70 0.88 0.62
3 clusters 9.00 3.50 1.90 0.85 0.60
5 clusters 8.40 3.70 1.90 1.10 0.72
Table 2: The optimal scaled densities ($d]{\delta}$) of KS-CCDs on synthetic data sets consisting of uniform clusters.
@@
Table 3: The simulation results of RK-CCDs on a set of synthetic data sets consisting of uniform clusters.
@@
Table 4: The simulation results of UN-CCDs on a set of synthetic data sets consisting of uniform clusters.

### v2

@@ 3 $v_{temp}\leftarrow\arg\max_{v\in V(D)}\{sc(v)\}$;
[…]
 6 $D\leftarrow D_{sub}(V_{temp})$;
@@
Greedy Algorithm 2 A greedy algorithm to find an approximate MDS which is greedy in a score function $sc(v)$ at each iteration (digraph is updated after each iteration). $D_{sub}(S)$ is the induced sub-digraph of vertex set $S$ from a digraph $D$.
2.2.1 Cluster Catch Digraphs Based on a KS-Type Statistic
@@
CCDs, introduced by Devinney [devinney2003class], are an adaptation of CCCDs [marchette2003classification] for unsupervised clustering. Given an unlabeled dataset $\mathrm{X}=\{x_{1},x_{2},...,x_{n}\}\subset\mathbb{R}^{d}$, where points are drawn from a mixture distribution, the objective is to determine the number of clusters and the optimal data partitioning. Unlike CCCDs, CCDs determine the radius of each covering ball $B(x_{i},r_{x_{i}})$ (centered at $x_{i}$ with radius $r_{x_{i}}$) using a Kolmogorov-Smirnov (KS)-type statistic, which measures the deviation of the local intensity around each point $x_{i}$ from a null model such as the Complete Spatial Randomness (CSR, where points are distributed uniformly within a given area) [marchette2005random], the KS-type statistic for a point $x_{i}$ is defined as:
$T_{KS}(x_{i},r)=F_{RW}(x_{i},r)-F_{0}(x_{i},r),$ (1)
@@
where $F_{RW}(x_{i},r)$ represents the number of points within the covering ball $B(x_{i},r)$, and $F_{0}(x_{i},r)$ is the expected number of points under the null model, typically given by $F_{0}(x_{i},r)=\delta r^{d}$ under the assumption of CSR, where $\delta$ is a density parameter and $d$ is the dimension of the space.
The radius $r_{x_{i}}$ of a covering ball $B(x_{i},r_{x_{i}})$ is chosen to maximize its KS-type statistic:
$r_{x_{i}}=\arg\max_{r\geq 0}\{T_{KS}(x_{i},r)\},$ (2)
@@
which ensures that the points covered by $B(x_{i},r_{x_{i}})$ (including $x_{i}$) form a significant local cluster.
@@
Once the optimal radii are determined for all covering balls, a digraph for $\mathrm{X}$ is constructed, denoted as $D=(V(D),A(D))$, where $V(D)=\mathrm{X}$, and there is an arc from $x_{i}$ to $x_{j}$ ($x_{i}x_{j}\in A(D)$) if $x_{j}\in B(x_{i},r_{x_{i}})$. Once the digraph $D$ is constructed, the weakly connected components of $D$ can be taken as clusters.
@@
In order to prevent the impact of noise, one may want to reduce the “complexity” of the covering balls. To achieve this goal, we apply the Greedy Algorithm 1 to find an approximate MDS, denoted as $\hat{S}$, which selects a subset of covering balls, representing the clusters more efficiently [marchette2005random]. The algorithm iteratively selects the vertex with the largest outdegree (i.e., the vertex with covering ball covering the most points) until all points are covered.
@@
To further reduce the complexity of the cluster cover, an intersection graph $G_{MD}=(V_{MD},E_{MD})$ is constructed, where $V_{MD}=\hat{S}$, and an edge exists between $u,v\in\hat{S}$ if their covering balls cover some common points, i.e., $\bar{N}(u)\cap\bar{N}(v)\neq\emptyset$. The Greedy Algorithm 2 is then applied to $G_{MD}$ to prune the MDS $\hat{S}$ further, yielding a reduced set $\hat{S}(G_{MD})$ that represents the center of latent clusters, and an initial partition $P=\{P_{1},P_{2},...,P_{k}\}$ for $\mathrm{X}$ can be constructed based on $\hat{S}(G_{MD})$.
Finally, to remove noise or outlying clusters, Manukyan and Ceyhan [manukyan2019parameter] designed an automatic refinement process based on silhouette index (a metric measures the quality of clustering) [gan2020data]. This process first ranks the partitions in $P$ in a decreasing order based on their size. Starting from the first two, it adds partitions incrementally as valid clusters until the average silhouette index of the entire dataset is maximized, indicating the optimal number of clusters is reached. In summary, the refinement process works on a set of candidate partitions (or clusters) and the final, selected ones are the finalized clusters, whose covering balls are the dominating covering balls.
During the above automatic refinement process, re-partitioning occurs after each incremental addition of a new partition, in order to re-evaluate the overall average silhouette index. In the re-partitioning step, points are assigned to the existing partition with the smallest relative dissimilarity measure $\rho(x_{i},B(x_{j},r_{x_{j}}))$, which is defined as follows,
$\rho(x_{i},B(x_{j},r_{x_{j}}))=\frac{d(x_{i},x_{j})}{r_{x_{j}}},$ (3)
where $d(x_{i},x_{j})$ is any distance or dissimilarity measure and $B(x_{j},r_{x_{j}})$ is the covering ball for a existing partition during this process.
2.2.2 Cluster Catch Digraphs Based on Ripley’s K Function
@@
Although the KS-CCD method is robust to noise clusters thanks to the employment of the silhouette index, it still has some limitations. For example, the KS-CCD method is not able to capture the spatial distributions of each cluster, and it need an input density parameter $\delta$ to determine the radius of each covering ball [manukyan2019parameter]. To address those limitations, instead of using the KS-type statistic, Manukyan and Ceyhan [manukyan2019parameter] applied the Ripley’s K function [ripley1976second], denoted as $K(t)$, to construct a test determining whether the points within each covering ball follow a homogeneous Poisson process (HPP, i.e., Complete Spatial Randomness). This test is referred to as the Spatial Randomness Monte Carlo Test (MC-SRT) with Ripley’s K function.
More specifically, given a dataset $\mathrm{X}$ consists of i.i.d points $\{x_{1},x_{2},...,x_{n}\}$ in a region $R\subset\mathbb{R}^{d}$, the Ripley’s $K$ function measures the spatial distribution of $\mathrm{X}$, revealing whether the points of $\mathrm{X}$ are clustered, dispersed, or randomly distributed. It is defined as follows [dixon2012ripley],
[…]
where $\lambda$ is the density of the point set.
$K$ function can be applied to describe certain point patterns of $\mathrm{X}$ and the most common application is to test CSR [dixon2012ripley], where the test statistic is the estimator of $K(t)$, denoted as $\widehat{K}(t)$, which is proportional to the ratio of the actual number of pairs whose distance is less than $t$ against the intensity of the point set. For more details, please refer to Manukyan and Ceyhan’s work [manukyan2019parameter].
@@
The RK-CCD method sets the radius of each covering ball to the largest value for which the points inside the ball fail to reject the null hypothesis of CSR, ensuring that the points encapsulated conform to a CSR pattern. The primary difference between RK-CCDs and KS-CCDs is in the determination of the radii, which is now based on Ripley’s K function rather than the KS-type statistic. Subsequent steps, such as reducing cover complexity using an approximate MDS, and refining the clustering result by maximizing the silhouette index, remains the same as in KS-CCDs.
@@
A variant of RK-CCDs introduced by Manukyan and Ceyhan allows for the detection of clusters with arbitrary shapes by considering the connected components of the intersection graph $G_{MD}$ as the final clusters instead of using the MDS [manukyan2019parameter].
@@ 3.1 The Limitations of RK-CCDs for Clustering
In our previous work,, we proposed the U-MCCD and SU-MCCD methods for outlier detection [shi2024outlier], where SU-MCCDs are the “flexible” versions of U-MCCD, adaptable for arbitrarily shaped clusters. As their first step, both methods employ RK-CCDs for clustering. However, the Monte Carlo analysis reveals that outlier detection with U-MCCDs and SU-MCCDs exhibit high False Positive Rates (TNRs) for regular points (falsely identifying regular points as outliers) in high-dimensional datasets. We have shown that this degradation is attributed to the limitations of RK-CCDs in high-dimensional spaces, whose dominating covering balls are not sufficiently large for latent clusters [shi2024outlier].
@@
Specifically, RK-CCDs present two key limitations [shi2024outlier]: (1) The initial points in the MC-SRT are not random, as they are always the centers of the covering balls, leading to inaccurate testing results, especially in high-dimensional datasets; (2) In MC-SRT, the point-wise confidence band for $K(t)$ is constructed on fixed $t$ values [manukyan2019parameter], which is not appropriate when the dimensions are not fixed, especially for high-dimensional datasets, where the distances between any two points get arbitrarily large.
3.2 The MC-SRT with the Nearest Neighbor Distance
@@
To avoid the limitations of RK-CCDs, we have introduced a new CCD approach for several new outlier detection methods and scores [shi2024outlier, shi2025outlierscores]. This new CCD conduct MC-SRT with the Nearest Neighbor Distance (NND) instead of the Ripley’s K function.
@@
Given a dataset $\mathrm{X}=\{x_{1},x_{2},...,x_{n}\}$, the mean NND of $\mathrm{X}$ is $\bar{\mathrm{d}}=\frac{\sum_{i=0}^{n}\mathrm{d}_{i}}{n}$, where $\mathrm{d}_{i}$ represents the distance of $x_{i}$ to its nearest neighbor. We aim to compares $\bar{\mathrm{d}}$ with the expected mean NND (denoted as $\mu_{\mathrm{d}}$) under the assumption of CSR. A normality test offers a convenient way to obtain $p$-values for $\bar{\mathrm{d}}$. Because both $\mu_{\mathrm{d}}$ and $\sigma_{\bar{\mathrm{d}}}$ (the standard derivation of $\bar{\mathrm{d}}$ under the assumption of CSR) can be derived [clark1954distance, clark1979generalization]. However, when the sample size $n$ is small, the distribution of $\bar{\mathrm{d}}$ is skewed, making the normality test unfeasible [besag1977simple]. Fortunately, a Monte Carlo test proposed by Besag and Diggle [besag1977simple] provides an alternative, where the $p$-values are determined by the empirical quantiles.
@@
Consequently, we have proposed another CCD approach based on the MC-SRT with NND. Specifically, we made several key enhancements to boost the robustness and reliability of the new CCDs [shi2024outlier]: (1) We conduct MC-SRT with both median and mean NND, applying a Holm’s Step-Down Procedure for correction [witten2013introduction], which improves the accuracy of the test substantially; (2) Our previous work shows that larger covering balls for each point are preferred for high-dimensional datasets [shi2024outlier]. Therefore, we offer the option to test the candidate radius values in a descending order, stops when the current radius value is not rejected in MC-SRT. (3) The MC-SRT is lower-tailed as the upper tail indicates the point pattern is significantly “regular”, which is less relevant to our interest. (4) The center points are excluded from the test as they are not random.
@@
The new clustering approach based on CCDs and NND is called Uniformity- and Neighbor-based CCD (UN-CCD) clustering method, which is presented in Algorithm 3. It is worth noting that the only difference among KS-CCDs, RK-CCDs, and UN-CCDs is how to determine the radius of covering balls. The time and space complexities of UN-CCDs for clustering are $O((N+d)n^{2}+n^{3})$ and $O(n^{2})$ respectively, which we have proven in our previous work [shi2024outlier].
We summarize the time and space complexity of RK-CCDs, KS-CCDs, and UN-CCDs in Table 1. When $N$ and $d$ are fixed, KS-CCDs and UN-CCDs have the time complexity of $O(n^{3})$, while RK-CCDs run in $O(n^{3}\log n)$, slower than the other two. All the three methods have the same space complexity of $O(n^{2})$.
Input: $\alpha$, dataset $\mathrm{X}=\{x_{1},x_{2},...,x_{n}\}$;
Output: Cluster centers of $\mathrm{X}$;
8pt Algorithm Steps:
1 foreach $x_{i}\in\mathrm{X}$ do
 2 Calculate distances $D(x_{i})=\{d(x_{i},x_{j})|x_{j}\in\mathrm{X},x_{i}\neq x_{j}\}$;
 35pt foreach distance $r_{(j)}$ in $D(x_{i})$ sorted ascending do
    4 Perform the MC-SRT with NND on the $\mathrm{X}\cap B(x_{i},r_{(j)})$;
    55pt if test rejected at level $\alpha$ then
@@ 6 Set $r_{x_{i}}=r_{(j-1)}$; break;
75pt Construct a CCD, $D=(V,A)$ using the pre-determined radii;
@@ 11 Output $\hat{S}(G_{MD})$ as cluster centers;
Algorithm 3 (UN-CCD Clustering method) Cluster Catch Digraphs based on the Nearest Neighbor Distance (NND). $\alpha$ is the level of the Monte Carlo test with NND.
Algorithms Time Complexity Space Complexity
KS-CCDs $O(n^{3}+n^{2}(d+\log n))$ $O(n^{2})$
RK-CCDs $O(n^{3}(\log n+N)+n^{2}d)$ $O(n^{2})$
UN-CCDs $O((N+d)n^{2}+n^{3})$ $O(n^{2})$
Table 1: The time/space complexity of all CCD-based methods.
@@
In Section 4, we compare the performance of KS-CCDs, RK-CCDs, and UN-CCDs with synthetic datasets consisting of multiple clusters. We show the advantage of UN-CCDs over RK-CCDs on Gaussian clusters and in high-dimensional settings.
4 Monte Carlo Experiments for Clustering Based on UN-CCDs
4.1 Experiment with Uniform Settings
@@
Similar to the design of simulations in our previous work [shi2024outlier], the analysis begins with datasets consisting of uniform clusters.
@@
Additionally, the cluster centers are:
  • $1^{st}$ cluster: $\bm{\mu_{1}}=(\underbrace{3,...,3}_{d})$;
  • $2^{nd}$ cluster: $\bm{\mu_{2}}=(6,\underbrace{3,...,3}_{d-1})$;
  • $3^{rd}$ cluster: $\bm{\mu_{3}}=(6,6,\underbrace{3,...,3}_{d-2})$;
  • $4^{th}$ cluster: $\bm{\mu_{4}}=(3,5.5,\underbrace{3,...,3}_{d-2})$;
  • $5^{th}$ cluster: $\bm{\mu_{5}}=(8.5,\underbrace{3,...,3}_{d-1})$.
The minimal distances between the first three cluster centers are 3, ensuring well separate clusters. On the other hand, the distances between the $1^{st}$ and the $4^{th}$, the $2^{nd}$ and the $5^{th}$ cluster centers are 2.5, leading in closer clusters, which may be confused as one cluster by many clustering methods. This study investigates the performance of the three CCD-based clustering methods for close clusters. We present three realizations with different number of clusters under 2-dimensional space in Figure 2.
(a) 2 clusters
(b) 3 clusters
(c) 5 clusters
@@
Clustering quality is assessed using both internal and external validation measures, including the Adjusted Rand Index (ARI) [hubert1985comparing], the average silhouette index (Sil) [rousseeuw1987silhouettes], and the frequency of successfully identifying the corrected number of clusters (denoted as success rate or SR for simplicity). We repeat each simulation 500 times, and the average measures across these repetitions are recorded.
To accommodate the decreased data intensity in high-dimensional datasets, one may increase the volumes of covering balls as dimensions increase. A smaller significance level $\alpha$ of MC-SRT (for both RK-CCDs and UN-CCDs) leads to larger radius for the covering balls. Therefore, we can achieve this objective by tuning the significance level $\alpha$ of MC-SRT (for both RK-CCDs and UN-CCDs), After some trial-and-error process (not shown here), we set $\alpha$ to $1\%$ for $d<10$ and $0.1\%$ for $d\geq 10$ for RK-CCDs, which achieves the best overall performance. For UN-CCDs, the $\alpha$ of the MC-SRT (with NND) is set to $\{15\%,10\%,5\%,1\%,0.1\%\}$ as $d$ increases from 2 to 20 for the optimal performance. Unlike RK-CCDs and UN-CCDs, KS-CCDs require a density parameter $\delta$ for the KS-type statistic. To manage the scale of $\delta$ in high-dimensional settings, We work on the value of $\sqrt[d]{\delta}$ instead, and select the $\sqrt[d]{\delta}$ maximizing the average silhouette index in each simulation, presented in Tables 2 and 6. The simulation results are summarized in Tables 3, 4, and 5. For better visualization, we summarize the simulation results in the following line plots (Figures 3 to 5), illustrating the trend of the performance (ARIs) of each CCD-based methods with varying data sizes ($n$), dimensions ($d$), and the number of clusters. To keep the figures uncluttered, we fix the number of clusters to 3 for the line plots with varying data sizes and dimensions; and the data size $n$ is fixed to 200 when we study the performance trend as the number of clusters increases.
Density Parameter: $\sqrt[d]{\delta}$ $d=2$ $d=3$ $d=5$ $d=10$ $d=20$
$n=50$ 2 clusters 3.70 2.00 1.30 0.75 0.49
3 clusters 3.10 1.80 1.20 0.700 0.55
5 clusters 2.60 1.60 1.10 0.800 0.69
$n=100$ 2 clusters 5.20 2.50 1.50 0.80 0.53
3 clusters 4.40 2.40 1.30 0.78 0.56
5 clusters 3.80 2.10 1.30 0.88 0.70
$n=200$ 2 clusters 7.00 3.00 1.50 0.85 0.60
3 clusters 5.90 3.10 1.50 0.82 0.57
5 clusters 5.40 2.70 1.50 0.98 0.74
$n=500$ 2 clusters 11.00 3.90 1.70 0.88 0.62
3 clusters 9.00 3.50 1.90 0.85 0.60
5 clusters 8.40 3.70 1.90 1.10 0.72
Table 2: The optimal scaled densities ($\sqrt[d]{\delta}$) of KS-CCDs on synthetic datasets consisting of uniform clusters.
@@
Table 3: The simulation results of KS-CCDs on a set of synthetic datasets consisting of uniform clusters. While KS-CCDs generally provide superior clustering performance, their effectiveness drops significantly when identifying a high number of clusters in a small dataset due to low cluster intensity.
Figure 3: The line plots of the ARIs of KS-CCDs, under the uniform cluster settings.
@@
Table 4: The simulation results of RK-CCDs on a set of synthetic datasets consisting of uniform clusters. The performance of RK-CCDs declines sharply in high-dimensional spaces due to the limitation of the Ripley’s $K$ function.
Figure 4: The line plots of the ARIs of RK-CCDs, under the uniform cluster settings.
@@
Table 5: The simulation results of UN-CCDs on a set of synthetic datasets consisting of uniform clusters. UN-CCDs perform comparably to RK-CCDs in low dimensions but are substantially superior in high dimensions.
Figure 5: The line plots of the ARIs of UN-CCDs, under the uniform cluster settings.